OpenStudy (anonymous):

(-3) is a solution of y=4-3x

OpenStudy (anonymous):

well, i assume here by solution you mean that -3=4-3x. Now if we subtract 4 from both sides we get: -7=-3x. Then we can divide both sides by -3 and get: 7/3=x.

OpenStudy (anonymous):

(-3 ) is a solution of y =4-3x

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

you did not show me where the y goes

OpenStudy (anonymous):

i simple replaced y with -3

OpenStudy (anonymous):

but i mean is -3 an x coordinate or y coordinate

OpenStudy (anonymous):

okay I see what you mean now

OpenStudy (anonymous):

you see what i mean, i am not sure if by solution you mean, that it could be a possible x value. Or do you mean that -3=4-3x and i am suppose to find the x

OpenStudy (anonymous):

-3 is the solution of y

OpenStudy (anonymous):

here is how it is writing -3___is a solution of y=4-3x.

OpenStudy (anonymous):

hello LagrangeSon678 are you there

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

what is suppose to go in the blank, do you have choices

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i mean what possibly completes that ?

OpenStudy (anonymous):

do you have answer?

OpenStudy (anonymous):

No it have got me puzzle

OpenStudy (anonymous):

but i mean you wrote: -3_____is a solution of y=4-3x? nothing can possibly go there

OpenStudy (anonymous):

yes that is the way it goes I will go with the frist answer you gave me

OpenStudy (anonymous):

thank you for your help

OpenStudy (anonymous):

no problem, come back if you figure it out, i want to know the answer now :)